On IQ distribution (normal) , with a mean of 100 and a standard deviation of 15 , find the actual IQs (raw scores)

for individuals with the following z scores.

1) .60

2) 2.60

3) -1.80

4) -.20

5) 2.80
Answer #1

\\ \mu = 100 \\\\ \sigma = 15 \\ \\ x = \mu + \sigma z = 100 + 15z \\ \\ x_1 = 100+15*0.6 = 109 \\\\x_2 = 100+15*2.6 = 139 \\\\x_3 = 100 + 15*(-1.8) = 73 \\\\x_4 = 100+15*(-0.2) = 97 \\\\x_5 = 100+15*2.8 = 142 \\\\$ Hope this was helpful. Please leave back any comment.$